# SPC: Evolving Self-Play Critic via Adversarial Games for LLM Reasoning

### v1

@@ 4.2 Main Results
Critic Performance on Reasoning Process Benchmarks   As shown in Tabs. 1 and 2, we compare our critic models with other baselines on 3 math-related reasoning process benchmarks to evaluate the abilities of predicting step correctness. We can observe that: (1) Our proposed SPC is gradually evolving and achieves state-of-the-art performance among all baselines. For example, the average performance on ProcessBench has improved from 70.8% to 77.7%, and on DeltaBench from 54.9% to 60.5%. (2) On all benchmarks, our method outperforms the latest PRMs specifically designed for scoring steps. (3) The performance of prompting LLMs as critics is not as good as SPC. Our method outperforms the distilled R1 model with the same size of 7B parameters. (4) Some baselines (PRMs and prompting Llama) have imbalanced recall between correct and error steps, leading to poor harmonic mean, whereas our critic is more balanced. (5) Our critic has not been trained with long CoT data but can successfully generalize, achieving the best performance on DeltaBench. In contrast, the two PRMs trained on short CoT show a significant decline in performance, with HarMean scores of only 14.3% and 41.3% on DeltaBench, respectively.
@@
The Effectiveness of Guiding Test-Time Search   Existing PRMs can enhance performance by ranking the completely generated reasoning steps or by aggregating scores using self-consistency [26, 2]. We apply the proposed SPC to LLM reasoning search, utilizing SPC to check the correctness of each step and regenerating the step if it is incorrect (up to 5 retries). Moreover, SPC can be combined with self-consistency by conducting a majority vote over several independent searches. For a fair comparison, all methods incorporating self-consistency sample 5 outputs in our experiments. In addition, for experiments without using self-consistency, we run them at least three times and average the results to reduce randomness. As shown in Tab. 3, on two popular benchmarks MATH500 [50] and AIME2024 [37], SPC significantly improves the performance of three types of LLM solvers, and outperforms five baseline verifiers. For instance, using the Qwen Solver at AIME2024, our SPC combined with Self-Consistency achieves a problem-solving accuracy of 23.3%, which is superior to the 16.7% accuracy of Self-Consistency + Qwen2.5-Math-7B-PRM800K. Notably, our SPC is trained using only short CoT data, yet it can still generalize to the DeepSeek-R1-Distill-Qwen-7B model, which outputs in a long CoT style. It achieved 94.0% accuracy on MATH500, whereas Math-Shepherd and Qwen2.5-Math-7B-PRM800K achieved only 89.2% and 91.8%, respectively.
@@
Table 2: Comparison of our SPC with baselines on the test set of PRM800K [23] and DeltaBench [35], where we extract human-annotated correct and erroneous steps for evaluating the correctness of step critiques. “Correct” and “Error” represent the recall of predicted critiques on correct and erroneous steps, respectively. “Average” denotes the arithmetic average of the recall of correct and erroneous steps, while “HarMean” represents their harmonic mean.
@@
Figure 3: Ablation study of our critic and sneaky generator. Left: the impact of different strategies on evolving critic models. Right: the success rate of sneaky generator attacking LLM solver and its win rate against the round 0 and round 1 critics.
@@ 4.3 Ablation Study
The Impact of Different Strategies on Evolving Critic   In Fig. 3 (left), we test critic models on ProcessBench, demonstrating the impact of different adversarial training methods. We refer to the sneaky generator and critic initialized after SFT as Sneaky-0 and Critic-0, respectively, while Sneaky-$n$ and Critic-$n$ represent models trained with $n$ rounds of self-play adversarial data. In round 1, Sneaky-1 and Critic-1 are trained using data generated from the adversarial game between Sneaky-0 and Critic-0. For each successfully transformed erroneous step, we have the critic predict four critiques, which may include both correct and incorrect predictions, forming a pair of positive and negative samples with the same input but different outputs. This method of constructing paired samples is more effective in RL training, improving the critic from 70.8% in round 0 to 76.8%, whereas not constructing paired samples only achieves a performance of 75.0%.
For round 2, we explore two adversarial settings. (1) Generating round 2 data using the confrontation between Sneaky-1 and Critic-1 and mixing it with the data from round 1. We observe a significant performance decline in the critic trained with this setting, dropping from 76.8% to 72.0%, possibly due to overfitting. We notice that the win rate of Sneaky-1 against Critic-1 is only 13.2%. Therefore, such an overly unbalanced game might prevent the critic from learning new knowledge from the adversarial process, similar to how humans need opponents of comparable skill levels when playing chess. Therefore, we adopt another setting: (2) Generating data through the game between Sneaky-1 and Critic-0, given that Sneaky-1 had a win rate of 32.5% against Critic-0. We then mix the data from both rounds for training Critic-0 and update it as Critic-2. Balancing the game prevents performance degradation and enables self-evolution, improving SPC’s performance to 77.7%.
The Performance of Sneaky Generator   As shown in Fig. 3 (right), we analyze sneaky generators’ success rates in attacking Qwen-2.5-7B-Instruct solver, as well as their win rates against Critic-0 and Critic-1. It is observed that the proportion of successful attacks on the solver gradually increases from 21.5% to 33.6%, as the sneaky generator iterates. We then feed successfully generated erroneous steps to the critic models. Sneaky generators’ win rates against Critic-0 increase from 20.6% (Sneaky-0) to 30.3% (Sneaky-2). Overall, the performance of the sneaky generators is iteratively improved.
@@ Appendix A Prompts
Appendix B More Details
B.1 Evaluation Details
@@ B.3 Training Hyperparameters
Appendix C Case Analysis
As shown in Fig. 9, we present a comparison of critiques provided by two SPCs trained in round 0 and round 2. The input to the SPCs includes not only the system prompt (Sec. A but also the problem, partial solution, and the last step of the current reasoning process within the blue box in the figure. The last step contains a logical error, mistakenly identifying two expressions as unmatched. The critique from the round 0 SPC considers the last step to be correct, agreeing with the view that the expressions are inconsistent. However, the round 2 SPC, evolved through self-play training, accurately identifies the type of error, namely a logical error (underlined in the figure), recognizing that the two sides of the equation can be equivalently substituted. The entire analysis process is clear and coherent, ultimately leading to the correct prediction that this step is incorrect.

### v2

@@ 4.2 Main Results
Critic Performance on Reasoning Process Benchmarks   As shown in Tabs. 1 and 2, we compare our critic models with other baselines on 3 math-related reasoning process benchmarks to evaluate the abilities of predicting step correctness. We can observe that: (1) Our proposed SPC is gradually evolving and achieves state-of-the-art performance among all models. For example, the average performance on ProcessBench has improved from 70.8% to 77.7%, and on DeltaBench from 54.9% to 60.5%. (2) On all benchmarks, our method outperforms the latest PRMs specifically designed for scoring steps. (3) The performance of prompting LLMs as critics is not as good as SPC. Our method outperforms the distilled R1 model with the same size of 7B parameters. (4) Some baselines (PRMs and prompting Llama) have imbalanced recall between correct and error steps, leading to poor harmonic mean, whereas our critic is more balanced. (5) Our critic is trained on short CoT data (from Qwen and Llama), but it is able to generalize to long CoT reasoning steps (e.g., R1 [21] and QwQ [22]) in DeltaBench. In contrast, the two PRMs trained on short CoT show a significant performance decline in DeltaBench, with HarMean scores of only 14.3% and 41.3%, respectively.
@@
The Effectiveness of Guiding Test-Time Search   Existing PRMs can enhance performance by ranking the completely generated reasoning steps or by aggregating scores using self-consistency [26, 2]. We apply the proposed SPC to LLM reasoning search, utilizing SPC to check the correctness of each step and regenerating the step if it is incorrect (up to 5 retries). Moreover, SPC can be combined with self-consistency by conducting a majority vote over several independent searches. For a fair comparison, all methods incorporating self-consistency sample 5 outputs in our experiments. In addition, for experiments without using self-consistency, we run them at least three times and average the results to reduce randomness. As shown in Tab. 3, on two popular benchmarks MATH500 [54] and AIME2024 [37], SPC significantly improves the performance of three types of LLM solvers, and outperforms five baseline verifiers. For instance, using the Qwen Solver at AIME2024, our SPC combined with Self-Consistency achieves a problem-solving accuracy of 23.3%, which is superior to the 16.7% accuracy of Self-Consistency + Qwen2.5-Math-7B-PRM800K. Notably, our SPC is trained using only short CoT data, yet it can still generalize to the DeepSeek-R1-Distill-Qwen-7B model, which outputs in a long CoT style. It achieves 94.0% accuracy on MATH500, whereas Math-Shepherd and Qwen2.5-Math-7B-PRM800K achieve only 89.2% and 91.8%, respectively.
@@
Table 2: Comparison of our SPC with baselines on the test set of PRM800K [23] and DeltaBench [35], where human-annotated correct and erroneous steps are extracted to evaluate the recall of critiques. “Correct” and “Error” represent the recall on correct and erroneous steps, respectively. “Average” denotes their arithmetic average and “HarMean” refers to their harmonic mean.
@@
Figure 3: Ablation study of our critic and sneaky generator. Left: The impact of different strategies on evolving critic models. Right: The success rate of sneaky generator in attacking LLM solver and its win rate against round 0 and round 1 critics.
@@ 4.3 Ablation Study
The Impact of Different Strategies on Evolving Critic
In Fig. 3 (left), we test critic models on ProcessBench, demonstrating the impact of different adversarial training methods. We refer to the sneaky generator and critic initialized after SFT as Sneaky-0 and Critic-0, respectively, while Sneaky-$n$ and Critic-$n$ represent models trained with $n$ rounds of self-play adversarial data. In round 1, Sneaky-1 and Critic-1 are trained using data generated from the adversarial game between Sneaky-0 and Critic-0. For each successfully transformed erroneous step, we have the critic predict four critiques, which may include both correct and incorrect predictions, forming a pair of positive and negative samples with the same input but different outputs. This method of constructing paired samples is more effective in RL training, improving the critic from 70.8% in round 0 to 76.8%, whereas not constructing paired samples only achieves a performance of 75.0%.
For round 2, we explore two evolving strategies. (1) Generating round 2 data using the confrontation between Sneaky-1 and Critic-1 and mixing it with the data from round 1. We observe a significant performance decline in the critic trained with this setting, dropping from 76.8% to 72.0%, possibly due to overfitting. We notice that the win rate of Sneaky-1 against Critic-1 is only 13.2%. Therefore, such an overly unbalanced game might prevent the critic from learning new knowledge from the adversarial process, similar to how humans need opponents of comparable skill levels when playing chess. Therefore, we adopt another setting: (2) Generating data through the game between Sneaky-1 and Critic-0, given that Sneaky-1 had a win rate of 32.5% against Critic-0. We then mix the data from both rounds for training Critic-0 and update it as Critic-2. Balancing the game prevents performance degradation and enables self-evolution, improving SPC’s performance to 77.7%.
The Performance of Sneaky Generator
As shown in Fig. 3 (right), we analyze sneaky generators’ success rates in attacking Qwen-2.5-7B-Instruct solver, as well as their win rates against Critic-0 and Critic-1. It is observed that the proportion of successful attacks on the solver gradually increases from 21.5% to 33.6%, as the sneaky generator iterates. We then feed successfully generated erroneous steps to the critic models. Sneaky generators’ win rates against Critic-0 increase from 20.6% (Sneaky-0) to 30.3% (Sneaky-2). Overall, the performance of the sneaky generators is iteratively improved.
@@ 5 Conclusion
Appendices
Appendix A Limitations and Societal Impact
Our SPC continuously generates data for reinforcement learning through adversarial games and has achieved impressive results. However, our current experiments are limited to several representative mathematical reasoning tasks. In future work, we plan to extend our approach to more general domains to further demonstrate the potential of the proposed framework.
@@ Appendix B Prompts
Appendix C More Details
C.1 Evaluation Details
@@ C.3 Training Hyperparameters
Appendix D Case Analysis
As shown in Fig. 9, we present a comparison of critiques provided by two SPCs trained in round 0 and round 2. The input to the SPCs includes not only the system prompt (Sec. B but also the problem, partial solution, and the last step of the current reasoning process within the blue box in the figure. The last step contains a logical error, mistakenly identifying two expressions as unmatched. The critique from the round 0 SPC considers the last step to be correct, agreeing with the view that the expressions are inconsistent. However, the round 2 SPC, evolved through self-play training, accurately identifies the type of error, namely a logical error (underlined in the figure), recognizing that the two sides of the equation can be equivalently substituted. The entire analysis process is clear and coherent, ultimately leading to the correct prediction that this step is incorrect.